OpenStudy (anonymous):

Kenny consumes a whole grain bagel and a 6 ounce orange juice for a total of 140 calories. Timmy consumes 2 whole grain bagels and 6 ounce orange juice for a total of 200 calories. How many more calories are in a 6 ounce orange juice than in a whole grain bagel? 20 40 60 68 80

OpenStudy (anonymous):

this one is simple really..we want to set up whats called a system of equations. since we have 2 variables x= cals in bagel and y=cals in juice then we need to equations to solve this. and we have these 2 equations. they are 1x+6y=140 and 2x +6y=200

OpenStudy (anonymous):

can you solve these

OpenStudy (anonymous):

yep. and solve for y, right?

OpenStudy (anonymous):

40/3

OpenStudy (anonymous):

well i would substitute y=(140-x)/6 into the other equation, solver for x. then knowing x you can easily plug it in to get y. then if your question asks for 6 bagels or whatever just multiply your x( totals cals in bagels( by 6